OpenStudy (anonymous):

How to solve this... given f(x+2)-f(x) = (6x+4)^2 and f(0)=-16. Find the value of f(5) ?

OpenStudy (anonymous):

any ideas ?

OpenStudy (kainui):

Consider that you have a way to find f(2) since you know what f(0) is.

OpenStudy (anonymous):

i got f(2) = 0

OpenStudy (kainui):

How?

OpenStudy (anonymous):

i put x=0, f(2) - f(0) = 4^2 f(2) - (-16) = 16 f(2) = 0 right ?

OpenStudy (anonymous):

what;s next ? :)

OpenStudy (anonymous):

i confused how going to f(5) ?

OpenStudy (watchmath):

Is there any mention that f(x) is a cubic function? even if it is and assuming that the problem is true for any function satisfying the given condition, then f(5)=564

OpenStudy (watchmath):

even if it is NOT I meant. Notice that \(f(x)=6x^3-6x^2-4x-16\) satisfying the given condition.

OpenStudy (anonymous):

hmmmm... didnt say if f(x) is a cubic function

OpenStudy (watchmath):

OK, I think I have an idea. But it needs an assumption that \(f\) is differentiable three times. Define \(h(x)=f(x+2)-f(x)\). Show that \(h'''(x)\) is constant. From there I think we can conclude that \(f(x)\) is a cubic function

OpenStudy (kainui):

What kind of magic did you do to come up with that cubic equation in the first place? I'm dumbfounded on two accounts of this problem lol.

OpenStudy (anonymous):

i get the new problem, how to differentiate it ? :)

OpenStudy (anonymous):

I don't think it is possible to find f(x) if x is odd number with the given information.

OpenStudy (agent0smith):

^ yeah we can find f(4) and f(6), but idk how to find f(5)...

OpenStudy (shubhamsrg):

i verified what @watchmath said,,the eqn fits in! how did you get that ?

OpenStudy (agent0smith):

At least using f(4) and f(6) we can get an estimate of where the value of f(5) should lie.

OpenStudy (kainui):

You could try plotting all the even points and then find a line that fits. I have honestly no clue how he got that either.

OpenStudy (kainui):

Where did this problem come from anyways?

OpenStudy (sirm3d):

if begins with the assumption that \(f(x)\) is a cubic polynomial, and then curve-fitting using 4 different points.

OpenStudy (anonymous):

this is my friend's problem... but i cant help him

OpenStudy (kainui):

What class is your friend in?

OpenStudy (anonymous):

grade 12

OpenStudy (kainui):

I feel like there's a simple way to solve this that we just didn't notice.

OpenStudy (anonymous):

i think so, but i dont know by what ?

OpenStudy (anonymous):

This a second order difference equation with only one initial condition so the general solution is \[ f(x)=-a (-1)^x+a+6 x^3-6 x^2-4 x-18 (-1)^x+2 \] Where a is a constant. f(5) =600 +2 a where a can be anything